OpenStudy (mlgpro):

The hypotenuse of a right triangle is 50 mm long. One leg of the right triangle is 30 mm long. What is the length of the other leg?

OpenStudy (lasttccasey):

For right triangles:\[a^2+b^2 = c^2\] where c i the hypotenuse.

OpenStudy (acxbox22):

30^2+b^2=50^2 900+b^2=2500 b^2=1600 take square root of both sides b=40 mm